# Beyond the Laplacian: Interpolated Spectral Augmentation for Graph Neural Networks

### v1

@@ 1 Introduction
Graph neural networks (GNNs) are foundational tools for modeling and learning from relational data. GNNs operate by passing, transforming, and aggregating messages between neighboring nodes iteratively, thereby encoding connectivity and feature information into expressive node representations that can capture complex patterns and relationships. The GNN approach has seen significant success, enabling advances in domains ranging from biology (stokes2020deep) to recommender systems (ying2018graph; wu2022graph1). The effectiveness of GNNs, however, often hinges on the availability of rich and informative node-level features. In many practical settings, node features can be unavailable (due to reasons such as privacy, missing data, etc), limited, or corrupted with noise (rossi2022unreasonable). Naïve application of GNN methods in these settings can lead to degraded performance (said2023enhanced).
@@
In these settings where node features are limited or unavailable, the most natural approach is to perform feature augmentation. A common heuristic is to augment the node features with one-hot encoding vectors or the all-ones vector. Such features do not take into account the underlying graph’s topology. A more informative approach is spectral augmentation, where spectral embeddings derived from the graph Laplacian matrix’s eigenvectors are used as node features (said2023enhanced). Indeed, it is well known from the manifold learning and spectral clustering literature that Laplacian spectral embeddings capture rich structural properties of the underlying graph (ng2001spectral; von2007tutorial; belkin2003laplacian). This approach of pre-computing spectral embeddings as node features is natural, as it supplies prior structural information that even shallow GNNs can directly exploit, mitigating the need for expensive and deep architectures that often suffer from issues such as oversmoothing (rusch2023survey). Such embeddings have also been successfully applied in GNNs beyond the missing/limited features setting, such as in graph transformers (kreuzer2021rethinking; dwivedi2020generalization) as a form of positional encoding. lim2022sign showed that such Laplacian embeddings, after accounting for sign and basis symmetries, are expressive and can improve performance of GNNs.
@@
Laplacian spectral embeddings are often treated as the default option in GNN feature augmentation (said2023enhanced; lim2022sign). On the other hand, a variety of graph matrices motivated by different applications have been considered in the spectral graph theory and network science literature (grindrod2018deformed; ou2016asymmetric; haemers2011universal). These alternative matrices can induce spectral embeddings that capture different kinds of graph connectivity information. For example, priebe2019two showed that Laplacian and adjacency spectral embeddings capture drastically different structures of graphs that has substantial implications for downstream tasks such as clustering.
@@
Motivated by the success of spectral feature augmentation in GNNs, especially in the limited/missing node features setting (said2023enhanced; lim2022sign), we ask whether considering alternative spectral embeddings can lead to improved performance in practice. To this end, we propose a simple yet expressive family of graph matrices that contains many well known graph matrices as special cases. We call the resulting spectral embeddings captured by this family the interpolated Laplacian Embeddings (ILEs). Using tools from spectral graph theory, we provide interpretations for the information captured by ILEs. Through experiments, we show that the choice of graph matrix used to compute spectral embeddings can impact downstream graph learning performance in the missing/limited node features setting. We discuss how to select relevant matrices and tuning parameters in practice. Our results suggest that practitioners many benefit from considering a broader range of spectral embeddings when augmenting node features in GNNs.
@@ 1.1 Relevant Literature
We refer to hamilton2018inductiverepresentationlearninglarge; wu2022graph for general background on graph neural networks. Commonly used graph neural network architectures that we adopt in our experiments include Graph Convolutional Networks (GCN) (kipf2017semisupervisedclassificationgraphconvolutional), Graph Isomorphism Networks (GIN) (xu2019powerfulgraphneuralnetworks) and GraphSAGE (hamilton2018inductiverepresentationlearninglarge).
The idea of using Laplacian embeddings as positional encodings to augment input features for graph neural networks has been explored in the literature (said2023enhanced; dwivedi2021graph). Such Laplacian positional encodings are also used in graph transformers (dwivedi2020generalization; kreuzer2021rethinking). lim2022sign proposes SignNet and BasisNet as architectural components that can be used to process eigenvector embeddings to ensure invariance to symmetries that arILE from sign flips and rotations (when repeated eigenvalues are present).
In addition to the Laplacian, adjacency spectral embeddings have been explored in the spectral clustering literature (sussman2012consistent; cape2019spectral). General families of graph matrices and operators have been considered in the spectral graph theory and network science literature. Notable examples include the Katz similarity matrix (ou2016asymmetric), deformed Laplacian (grindrod2018deformed), the signless Laplacian (cvetkovic2007signless) and related generalizations (nikiforov2017merging), the universal Adjacency matrices (haemers2011universal), and many more. Different spectral embeddings can capture different graph structures with different interpretation, an observation noted in priebe2019two in the context of spectral clustering.
@@
In this work, we revisit the common practice of using Laplacian spectral embeddings for feature augmentation in graph neural networks. Our contributions are threefold: (i) we introduce interpolated Laplacian Embeddings (ILEs), which are eigenvectors derived from a simple and flexible family of graph matrices that generalizes many classical graph matrices as special cases; (ii) we provide theoretical analysis, leveraging tools from spectral graph theory, to characterize and interpret the structural information that ILEs encode; and (iii) we conduct extensive experiments demonstrating that the choice of graph matrix can lead to improvements in downstream GNN performance, especially under settings with limited or missing node features. Our results move beyond the default use of Laplacian embeddings, broadening the spectral augmentation toolkit available for GNNs.
@@ 2.1 Setup and Notation
We work in the setting of simple, undirected, connected, weighted graphs. Let $G=(V,E,w)$ denote such a graph, where $V$ is the vertex set, $E$ the edge set, and $w:E\to\mathbb{R}_{\geq 0}$ assigns a nonnegative weight to each edge. For convenience, we write $w_{e}$ or $w_{uv}$ to denote the weight of edge $e$ or edge $(u,v)\in E$. Throughout, we index vertices by $n]:=\{1,2,\dots,n\}$. We use $i,j,u,v$ to index vertices in $[n]$. We use the terms nodes and vertices interchangeably. We use bold fonts for matrices and vectors.
@@ 2.2 Spectral Embeddings of Nodes and The Two-Truths Phenomenon
The eigenvectors of graph matrices can be used to construct low-dimensional spectral node embeddings that preserves structural properties of the underlying graph. Consider the eigendecomposition of the Laplacian ${\mathbf{L}}=\sum_{u=1}^{n}\lambda_{u}{\mathbf{z}}_{u}{\mathbf{z}}_{u}^{T}$. A spectral embedding of dimension $k$ is constructed by selecting the first $k$ eigenvectors of $\mathbf{L}$} that corresponds to the smallest $k$ non-zero eigenvalues. The entries of such eigenvectors are then used as the coordinates in a Euclidean representation. More concretely, the embedding of node $u$ is given by the $u$-th row of ${\mathbf{z}}_{1},{\mathbf{z}}_{2},\dots,{\mathbf{z}}_{k}]\in\mathbb{R}^{n\times k}$. Adjacency spectral embeddings are constructed analogously, except the eigenvectors corresponding to the largest $k$ eigenvalues are selected.
Arguably the most canonical and commonly used graph matrices are the Laplacian matrix $\mathbf{L}$} and adjacency matrix $\mathbf{A}$}. Most work in the literature utilizes only one particular type of spectral embedding, usually the Laplacian by default. The effect of matrix choice on downstream tasks such as clustering have only been explicitly studied recently. Notably, the pioneering work of priebe2019two identified the “two-truths” phenomenon: spectral embeddings from the normalized Laplacian versus the adjacency matrix yield clustering results that are drastically different yet both valid. Laplacian spectral embeddings capture community structure, which reflects high connectivity within communities and low connectivity between communities. This can be interpreted via the well known relationship between Laplacian eigenvectors and approximate graph cuts (von2007tutorial). In contrast, adjacency embeddings capture “core-periphery” structure (priebe2019two). Here, high-degree nodes that are highly connected form a hub, whereas low-degree nodes that are sparsely connected (often only to the core) form the periphery. We pictorially illustrate this phenomenon in Figure 2(a). While priebe2019two used a stochastic blockmodel and asymptotic analysis to explain this phenomenon, in section 3.1 we provide an alternative explanation that does not resort to asymptotics. This phenomenon suggests that the choice of graph matrix for spectral embedding can affect downstream tasks, motivating our study of its impact on node classification with GNNs.
@@ 2.3 Graph Neural Networks and Spectral Embeddings
Graph neural networks (GNNs) most commonly operate under a message passing scheme, which iteratively applies two general steps: a message-passing step, where nodes aggregate information from their neighbors, and an update step, where nodes update their representations/learned embeddings using that aggregated information. The starting point of such representations is usually taken to be a set of node (and/or edge) features. Stacking $L$ such layers of message passing together yields node representation that encode structural information from neighborhoods up to $L$ hops away. This message passing scheme encompasses a wide variety of popular GNN architectures, including Graph Convolutional Networks (GCNs) (kipf2017semisupervisedclassificationgraphconvolutional), GraphSAGE (hamilton2018inductiverepresentationlearninglarge), and Graph Isomorphism Networks (GINs) (xu2019powerfulgraphneuralnetworks). The choice of how the messages are passed/aggregated and how the node representations are often what distinguishes GNN models from one another.
@@
Given the importance of node features for GNN learning, in many applications it is often of interest to supplement node features by spectral embeddings (said2023enhanced; lim2022sign). Since spectral embeddings are eigenvectors of symmetric matrices, they are subject to certain symmetries: for any eigenvector ${\mathbf{z}}\in\mathbb{R}^{n}$, the negated vector $\mathbf{z}$} is also an eigenvector associated with the same eigenvalue. Moreover, in cases of eigenvalue multiplicity, the corresponding eigenspace admits infinitely many valid orthonormal bases up to rotations. SignNet and BasisNet (lim2022sign) are general frameworks that allow for the use of spectral embeddings in GNNs while respecting symmetry constraints. SignNet achieves invariance to sign flips by symmetrizing the response of a neural network $\phi$ applied to each eigenvector individually. Given $k$ eigenvectors ${\mathbf{z}}_{1},\dots,{\mathbf{z}}_{k}\in\mathbb{R}^{n}$, SignNet defines the mapping $f({\mathbf{z}}_{1},\dots,{\mathbf{z}}_{k})\;=\;\rho\Big(\big[\phi({\mathbf{z}}_{i})+\phi(-{\mathbf{z}}_{i})\big]_{i=1}^{k}\Big),$ where $\rho$ is an additional learnable aggregation function. By construction, $f$ remains unchanged under any sign flip, thereby ensuring consistency of learned representations across different eigenbasis choices. BasisNet operates analogously for rotations for graphs that have repeated eigenvalues: given a matrix of eigenvectors ${\mathbf{Z}}=[{\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{k}]\in\mathbb{R}^{n\times k}$ as input, BasisNet considers the associated orthogonal projector matrix ${\mathbf{Z}}{\mathbf{Z}}^{\top}$ since it is invariant to rotations. We defer to lim2022sign for details. Rather than appending spectral embeddings directly as node features, SignNet and/or BasisNet first process the embeddings, and their outputs are then appended as features. While BasisNet is only applicable for eigenvectors with repeated eigenvalues, SignNet should always be used since sign-flips symmetry affects all eigenvectors.
@@ 3.1 A Family of Graph Matrices
where $\mathbf{J}$} is the all-ones matrix, $\mathbf{I}$} is the identity matrix, and $\alpha,\beta,\kappa,\zeta\in\mathbb{R}$. This family has natural mathematical properties and includes many of the most important graph matrices as special cases, such as the Laplacian, the signless Laplacian and the Seidel matrix (haemers2011universal). Despite the richness, the fact that this family depends on four tuning parameters makes it less practical for use in machine learning applications.
@@
where $t,s\in\mathbb{R}$. Here, we drop the $\mathbf{J}$} and $\mathbf{I}$} components from the universal adjacency matrix family, and without loss of generality adopt a subtraction based formulation to mimic the form of the Laplacian. We call spectral embeddings constructed from the eigenvectors of the interpolated Laplacian family of matrices interpolated Laplacian embeddings (ILEs). Below, we interpret the information captured by ILEs.
@@ 3.2 Interpretation of Interpolated Laplacian Embeddings
Thus, we can interpret the first $k$ eigenvectors of $\mathbf{M}$} as a set of vectors that collectively minimizes Rayleigh quotient. This optimization perspective provides a way for us to interpret the resulting embeddings.
@@
Consider the matrix family ${\mathbf{M}}(t,s)=t{\mathbf{D}}-s{\mathbf{A}},\qquad t,s\in\mathbb{R}$. For any unit vector ${\mathbf{x}}\in\mathbb{R}^{n}$, its quadratic form is given by

### v2

@@ 1 Introduction
Graph neural networks (GNNs) are foundational tools for modeling and learning from relational data. GNNs operate by passing, transforming, and aggregating messages between neighboring nodes iteratively, thereby encoding connectivity and feature information into expressive node representations that can capture complex patterns and relationships. The GNN approach has seen significant success, enabling advances in domains ranging from biology (stokes2020deep) to recommender systems (ying2018graph; wu2022graph1). The effectiveness of GNNs, however, often hinges on the availability of rich and informative node-level features. In many practical settings, node features can be unavailable (due to reasons such as privacy, missing data, etc.), limited, or corrupted with noise (rossi2022unreasonable). Naïve application of GNN methods in these settings can lead to degraded performance (said2023enhanced).
@@
In these settings where node features are limited or unavailable, the most natural approach is to perform feature augmentation. A common heuristic is to augment the node features with one-hot encoding vectors or the all-ones vector. Such features do not take into account the underlying graph’s topology. A more informative approach is spectral augmentation, where spectral embeddings derived from the graph Laplacian matrix’s eigenvectors are used as node features (said2023enhanced). Indeed, it is well known from the manifold learning and spectral clustering literature that Laplacian spectral embeddings capture rich structural properties of the underlying graph (ng2001spectral; von2007tutorial; belkin2003laplacian). This approach of pre-computing spectral embeddings as node features is natural, as it supplies prior structural information that even shallow GNNs can directly exploit, mitigating the need for expensive and deep architectures that often suffer from issues such as oversmoothing (rusch2023survey). Such embeddings have also been successfully applied in GNNs beyond the missing/limited features setting, such as in graph transformers (kreuzer2021rethinking; dwivedi2020generalization) as a form of positional encoding. lim2022sign showed that such Laplacian embeddings, after accounting for sign and basis symmetries, are expressive and can improve the performance of GNNs.
@@
Laplacian spectral embeddings are often treated as the default option in GNN feature augmentation (said2023enhanced; lim2022sign). On the other hand, a variety of graph matrices motivated by different applications have been considered in the spectral graph theory and network science literature (grindrod2018deformed; ou2016asymmetric; haemers2011universal). These alternative matrices can induce spectral embeddings that capture different kinds of graph connectivity information. For example, priebe2019two showed that Laplacian and adjacency spectral embeddings capture drastically different structures of graphs that have substantial implications for downstream tasks such as clustering.
@@
Motivated by the success of spectral feature augmentation in GNNs, especially in the limited/missing node features setting (said2023enhanced; lim2022sign), we ask whether considering alternative spectral embeddings can lead to improved performance in practice. To this end, we propose a simple yet expressive family of graph matrices that contains many well-known graph matrices as special cases. We call the resulting spectral embeddings captured by this family the Interpolated Laplacian Embeddings (ILEs). Using tools from spectral graph theory, we provide interpretations for the information captured by ILEs. Through experiments, we show that the choice of graph matrix used to compute spectral embeddings can impact downstream graph learning performance in the missing and limited node features settings. We discuss how to select relevant matrices and tuning parameters in practice. Our results suggest that practitioners may benefit from considering a broader range of spectral embeddings when augmenting node features in GNNs.
@@ 1.1 Relevant Literature
We refer to hamilton2018inductiverepresentationlearninglarge; wu2022graph for general background on GNNs. Commonly used GNN architectures that we adopt in our experiments include Graph Convolutional Networks (GCN) (kipf2017semisupervisedclassificationgraphconvolutional), Graph Isomorphism Networks (GIN) (xu2019powerfulgraphneuralnetworks), and GraphSAGE (hamilton2018inductiverepresentationlearninglarge).
The idea of using Laplacian embeddings as positional encodings to augment input features for graph neural networks has been explored in the literature (said2023enhanced; dwivedi2021graph). Such Laplacian positional encodings are also used in graph transformers (dwivedi2020generalization; kreuzer2021rethinking). lim2022sign proposes SignNet and BasisNet as architectural components that can be used to process eigenvector embeddings to ensure invariance to symmetries that arise from sign flips and rotations (when repeated eigenvalues are present).
In addition to the Laplacian, adjacency spectral embeddings have been explored in the spectral clustering literature (sussman2012consistent; cape2019spectral). General families of graph matrices and operators have been considered in the spectral graph theory and network science literature. Notable examples include the Katz similarity matrix (ou2016asymmetric), the deformed Laplacian (grindrod2018deformed), the signless Laplacian (cvetkovic2007signless) and related generalizations (nikiforov2017merging), the universal Adjacency matrices (haemers2011universal), and many more. Different spectral embeddings can capture different graph structures with different interpretations, an observation noted in priebe2019two in the context of spectral clustering.
@@
In this work, we revisit the common practice of using Laplacian spectral embeddings for feature augmentation in GNNs. Our contributions are threefold: (i) we introduce interpolated Laplacian Embeddings (ILEs), which are eigenvectors derived from a simple and flexible family of graph matrices that generalizes many classical graph matrices as special cases; (ii) we provide theoretical analysis, leveraging tools from spectral graph theory, to characterize and interpret the structural information that ILEs encode; and (iii) we conduct extensive experiments demonstrating that the choice of graph matrix can lead to improvements in downstream GNN performance, especially under settings with limited or missing node features. Our results move beyond the default use of Laplacian embeddings, broadening the spectral augmentation toolkit available for GNNs.
@@ 2.1 Setup and Notation
We work in the setting of simple, undirected, connected, and weighted graphs. Let $G=(V,E,w)$ denote such a graph, where $V$ is the vertex set, $E$ the edge set, and $w:E\to\mathbb{R}_{\geq 0}$ assigns a nonnegative weight to each edge. For convenience, we write $w_{e}$ or $w_{uv}$ to denote the weight of edge $e$ or edge $(u,v)\in E$. Throughout, we index vertices by $n]:=\{1,2,\dots,n\}$. We use $i,j,u,v$ to index vertices in $[n]$. We use the terms nodes and vertices interchangeably. We use bold fonts for matrices and vectors.
@@ 2.2 Spectral Embeddings of Nodes and The Two-Truths Phenomenon
The eigenvectors of graph matrices can be used to construct low-dimensional spectral node embeddings that preserve structural properties of the underlying graph. Consider the eigendecomposition of the Laplacian ${\mathbf{L}}=\sum_{u=1}^{n}\lambda_{u}{\mathbf{z}}_{u}{\mathbf{z}}_{u}^{T}$. A spectral embedding of dimension $k$ is constructed by selecting the first $k$ eigenvectors of $\mathbf{L}$} that correspond to the smallest $k$ non-zero eigenvalues. The entries of such eigenvectors are then used as the coordinates in a Euclidean representation. More concretely, the embedding of node $u$ is given by the $u$-th row of ${\mathbf{z}}_{1},{\mathbf{z}}_{2},\dots,{\mathbf{z}}_{k}]\in\mathbb{R}^{n\times k}$. Adjacency spectral embeddings are constructed analogously, except the eigenvectors corresponding to the largest $k$ eigenvalues are selected.
Arguably, the most canonical and commonly used graph matrices are the Laplacian matrix $\mathbf{L}$} and adjacency matrix $\mathbf{A}$}. Most work in the literature utilizes only one particular type of spectral embedding, usually the Laplacian by default. The effect of matrix choice on downstream tasks, such as clustering, has only been explicitly studied recently. Notably, the pioneering work of priebe2019two identified the “two-truths” phenomenon: spectral embeddings from the normalized Laplacian versus the adjacency matrix yield clustering results that are drastically different yet both valid. Laplacian spectral embeddings capture community structure, which reflects high connectivity within communities and low connectivity between communities. This can be interpreted via the well-known relationship between Laplacian eigenvectors and approximate graph cuts (von2007tutorial). In contrast, adjacency embeddings capture “core-periphery” structure (priebe2019two). Here, high-degree nodes that are highly connected form a hub, whereas low-degree nodes that are sparsely connected (often only to the core) form the periphery. We pictorially illustrate this phenomenon in Figure 2(a). While priebe2019two used a stochastic block model and asymptotic analysis to explain this phenomenon, in Section 3.1 we provide an alternative explanation that does not resort to asymptotics. This phenomenon suggests that the choice of graph matrix for spectral embedding can affect downstream tasks, motivating our study of its impact on node classification with GNNs.
@@ 2.3 Graph Neural Networks and Spectral Embeddings
GNNs most commonly operate under a message-passing scheme, which iteratively applies two general steps: a message-passing step, where nodes aggregate information from their neighbors, and an update step, where nodes update their representations or learned embeddings using that aggregated information. The starting point of such representations is usually taken to be a set of node (and/or edge) features. Stacking $L$ such layers of message passing together yields node representations that encode structural information from neighborhoods up to $L$ hops away. This message-passing scheme encompasses a wide variety of popular GNN architectures, including Graph Convolutional Networks (GCNs) (kipf2017semisupervisedclassificationgraphconvolutional), GraphSAGE (hamilton2018inductiverepresentationlearninglarge), and Graph Isomorphism Networks (GINs) (xu2019powerfulgraphneuralnetworks). The choice of how the messages are passed or aggregated and how the node representations are updated is often what distinguishes GNN models from one another.
@@
Given the importance of node features for GNN learning, in many applications it is often of interest to supplement node features by spectral embeddings (said2023enhanced; lim2022sign). Since spectral embeddings are eigenvectors of symmetric matrices, they are subject to certain symmetries: for any eigenvector ${\mathbf{z}}\in\mathbb{R}^{n}$, the negated vector $\mathbf{z}$} is also an eigenvector associated with the same eigenvalue. Moreover, in cases of eigenvalue multiplicity, the corresponding eigenspace admits infinitely many valid orthonormal bases up to rotations. SignNet and BasisNet (lim2022sign) are general frameworks that allow for the use of spectral embeddings in GNNs while respecting symmetry constraints. SignNet achieves invariance to sign flips by symmetrizing the response of a neural network $\phi$ applied to each eigenvector individually. Given $k$ eigenvectors ${\mathbf{z}}_{1},\dots,{\mathbf{z}}_{k}\in\mathbb{R}^{n}$, SignNet defines the mapping $f({\mathbf{z}}_{1},\dots,{\mathbf{z}}_{k})\;=\;\rho\Big(\big[\phi({\mathbf{z}}_{i})+\phi(-{\mathbf{z}}_{i})\big]_{i=1}^{k}\Big),$ where $\rho$ is an additional learnable aggregation function. By construction, $f$ remains unchanged under any sign flip, thereby ensuring consistency of learned representations across different eigenbasis choices. BasisNet operates analogously for rotations for graphs that have repeated eigenvalues: given a matrix of eigenvectors ${\mathbf{Z}}=[{\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{k}]\in\mathbb{R}^{n\times k}$ as input, BasisNet considers the associated orthogonal projector matrix ${\mathbf{Z}}{\mathbf{Z}}^{\top}$ since it is invariant to rotations. We defer to lim2022sign for details. Rather than appending spectral embeddings directly as node features, SignNet and/or BasisNet first process the embeddings, and their outputs are then appended as features. While BasisNet is only applicable for eigenvectors with repeated eigenvalues, SignNet should always be used since sign-flip symmetry affects all eigenvectors.
@@ 3.1 A Family of Graph Matrices
where $\mathbf{J}$} is the all-ones matrix, $\mathbf{I}$} is the identity matrix, and $\alpha,\beta,\kappa,\zeta\in\mathbb{R}$. This family has natural mathematical properties and includes many of the most important graph matrices as special cases, such as the Laplacian, the signless Laplacian, and the Seidel matrix (haemers2011universal). Despite the richness, the fact that this family depends on four tuning parameters makes it less practical for use in machine learning applications.
@@
where $t,s\in\mathbb{R}$. Here, we drop the $\mathbf{J}$} and $\mathbf{I}$} components from the universal adjacency matrix family, and without loss of generality adopt a subtraction-based formulation to mimic the form of the Laplacian. We call spectral embeddings constructed from the eigenvectors of the interpolated Laplacian family of matrices Interpolated Laplacian Embeddings (ILEs). Below, we interpret the information captured by ILEs.
@@ 3.2 Interpretation of Interpolated Laplacian Embeddings
Thus, we can interpret the first $k$ eigenvectors of $\mathbf{M}$} as a set of vectors that collectively minimizes the Rayleigh quotient. This optimization perspective provides a way for us to interpret the resulting embeddings.
@@
Consider the matrix family ${\mathbf{M}}(t,s)=t{\mathbf{D}}-s{\mathbf{A}}$, where $t,s\in\mathbb{R}$. For any unit vector ${\mathbf{x}}\in\mathbb{R}^{n}$, its quadratic form is given by